A sample of 25 women had a variance in IQ scores of 62. A sample of 18 men had a variance of 72. Do the women have a smaller variance in IQ scores at the 0.01 level of significance? Assume both populations are normally distributed


1
Expert's answer
2022-05-11T09:23:21-0400

The following null and alternative hypotheses need to be tested:

"H_0:\sigma_1^2\ge\sigma_2^2"

"H_a:\sigma_1^2<\sigma_2^2"

This corresponds to a left-tailed test, for which a F-test for two population variances needs to be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df_1=n_1-1=24" degrees of freedom, "df_2=n_2-1=17" degrees of freedom, and the rejection region for this left-tailed test test is "R = \{F: F < 0.3548\}."

The F-statistic is computed as follows:


"F=\dfrac{s_1^2}{s_2^2}=\dfrac{62}{72}=0.8611"

Since from the sample information we get that "F = 0.8611 \ge0.3548= F_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population variance "\sigma_1^2" is less than the population variance "\sigma_2^2," at the "\alpha = 0.01" significance level.